4) Suppose a random variable X has theprobability distribution with a: o 1 -2 0 1 2 0.3 0.1 p 0.4 . then p - ,P(X2 22) = ,, and E(X) = - 5) Suppose X~Bin(10,0.4), Y-2X+5, then E(Y) = ,Var(Y) 6) Suppose X-NC-3,4) and Y~N(2,9), X and Y are independent, then Var(X-2Y)
Answer #1

4)

p = 1 - (0.3 + 0.1 + 0.4) = 0.2

Distribution of X2 is

P(X2 = 4) = 0.3 + 0.4 = 0.7

P(X2 = 0) = 0.1

P(X2 = 1) = 0.2

So, P(X2\ge 2) = 0.7

E(X) = -2 * 0.3 + 0 * 0.1 + 1 * 0.2 + 2 * 0.4 = 0.4

5)

E[X] = 10 * 0.4 = 4

Var[X] = 10 * 0.4 * (1-0.4) = 2.4

E(Y) = E(2X + 5) = 2E(X) + 5 = 2 * 4 + 5 = 13

Var(Y) = Var(2X + 5) = 22Var(X) + 0 = 4 * 2.4 = 9.6

6)

If X and Y are independent,

Var(X - 2Y) = Var(X) + (-2)2 Var(Y) = 4 + 4 * 9 = 40